OpenStudy (anonymous):

Solve Algebra Equation: 5x+10=(x)/(3)=24. i kno the answer eqauls 3 but can someone explain me the steps???

OpenStudy (anonymous):

you have something wrong in the posted equation x/3 = 24 x=72 5x=10=24 5x=14 x=14/5 obviously a problem

OpenStudy (anonymous):

its 5x+10-x/3=24

OpenStudy (anonymous):

no u guys its x over 3

OpenStudy (saifoo.khan):

the question in incorrect, there are 2 equal signs.

OpenStudy (anonymous):

OOOOHHHH TYPO

OpenStudy (anonymous):

soory lemme fix it one sec

OpenStudy (saifoo.khan):

1 sec up!

OpenStudy (anonymous):

lol hold on

OpenStudy (saifoo.khan):

Holding..

OpenStudy (anonymous):

5x+10=(x)/(3)+24

OpenStudy (anonymous):

relax mister.... you will get stress

OpenStudy (anonymous):

MISTER IM A GIRL

OpenStudy (saifoo.khan):

lol @ sat.

OpenStudy (anonymous):

dnt make fun of me im jus a little girl :(

OpenStudy (anonymous):

i was talking to saifoo

OpenStudy (anonymous):

oops my bad *blushes

OpenStudy (saifoo.khan):

5x+10=(x)/(3)+24 \[5x + 10 = \frac{x + 72}{3}\] Cross multiply.. \[3(5x + 10) = x+72\]

OpenStudy (saifoo.khan):

Can you solve it now?

OpenStudy (anonymous):

he was being impatient. your equation is \[5x+1 =\frac{x}{3}+24\] right?

OpenStudy (anonymous):

oh yea i get i t im sooo dumb :P thnx soo much <3

OpenStudy (saifoo.khan):

You welcome!!! :)

OpenStudy (anonymous):

@ satelite no its 10 not 1 but saifoo answered it already but thnx alot to u both :)

OpenStudy (anonymous):

5x+10-x/3=24

OpenStudy (anonymous):

What they are really asking you is FIND the value of X

OpenStudy (anonymous):

lets reorder our equation by putting all the number values on to the right of the equal simbol

OpenStudy (anonymous):

5x-x=24-10(3) remember you can "move" anything from left to right if you follow the rules as: anything multipliying WILL pass to the other side Dividing and anything with the simbol of plus WILL pas to the other side with minus AND VICEVERSA

OpenStudy (anonymous):

so 5x-x is 4x right!! and 24-30 "because 10 times 3"